OpenStudy (anonymous):

The perimeter of a retangler is 194 inches. The length exceeds the width by 29 inches. Find the width and the lenght please

OpenStudy (anonymous):

You are nirvana. Am I right?

OpenStudy (anonymous):

et width=x let length =x+29 so zer are 2 length n 2 width 2(x)+2(x+29)=194 2x+2x+58=194 x=34 that is width=34 now length=29+34=63 zats it

OpenStudy (anonymous):

You are nirvana. Am I right?

OpenStudy (anonymous):

nirvana

OpenStudy (anonymous):

You are related with that term, somehow. Maybe thats your name or something

OpenStudy (anonymous):

Thanks again...what the heck does that mean? both right

OpenStudy (anonymous):

You know what that means....

OpenStudy (anonymous):

Anyhow, no more help from me...

OpenStudy (anonymous):

If you want you may take back those medals

OpenStudy (anonymous):

ok thanks I am not looking for you just from you Pi just needing help to find the answers to the last few problems. But thanks again.